• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name H. Benaoum

    Log entry time 00:11:16 on July31,2005

    Entry number 150150

    keyword=Swing Shift Summary


    16:00 Start Shift with an access to the Hall because of lost communication with ROC26 in the right arm.
    16:09 start run 3974 .
    16:15 problem fixed and we asked for 35uA.
    16:40 called MCC to ask for increase in current to 38uA.
    17:41 IOC target reboot (lost communication) .
    19:35 Seen that the right upper coil for Septa was increasing.It reaches at some point 7.642K... I asked MCC to lower the beam current to 35uA.
    19:36 Temperature of the right upper looks fine now 7.626K.
    23;30 called MCC for beam off because of CHL went down to 17 .
    23:50 MCC has RF problem ....
    24:05 MCC said will has beam back soon .

    Production runs :
    3974 IHWP OUT 100k
    3975 IHWP OUT 100k
    3976 junk
    3977 IHWP IN 100k
    3978 IHWP IN 100k
    3979 IHWP IN 100k
    3980 IHWP IN 100k
    3981 IHWP IN 100k
    3982 IHWP IN 100k
    3983 IHWP IN still running passed to owl shift